# Aircraft Trajectory Dataset Augmentation in Latent Space

### v1

@@ Abstract
Aircraft trajectory modeling plays a crucial role in Air Traffic Management (ATM) and is important for various downstream tasks, including conflict detection and landing time prediction. Dataset augmentation through the addition of synthetically generated trajectory data is necessary to develop a more robust aircraft trajectory model and ensure that the trajectory dataset is sufficient and balanced. In this work, we propose a novel framework called ATRADA for aircraft trajectory dataset augmentation. In the proposed framework, a Transformer encoder learns the underlying patterns in the original trajectory dataset and converts each data point into a context vector in the learned latent space. The converted dataset in the latent space is projected into reduced dimensions using principal component analysis (PCA), and a Gaussian mixture model (GMM) is applied to fit the probability distribution of the data points in the reduced-dimensional space. Finally, new samples are drawn from the fitted GMM, the dimension of the samples is reverted to the original dimension, and they are decoded with a Multi-Layer Perceptron (MLP). Several experiments demonstrate that the framework effectively generates new, high-quality synthetic aircraft trajectory data, which were compared to the results of several baselines.
Machine Learning, ICML
@@ 1 Introduction
As the demand for air transportation continues to increase around the world, the concept of air traffic operations has shifted to Trajectory-Based Operation (TBO). TBO more accurately manages an aircraft’s four-dimensional trajectory (4DT) in latitude, longitude, altitude, and time. Many studies have developed data-driven aircraft trajectory models for various downstream applications such as conflict detection and landing time prediction (Liu & Hwang, 2011; Hong & Lee, 2015).
One study on trajectory modeling suggested using a GMM to learn statistical patterns of aircraft trajectories (Barratt et al., 2018), while another study used a Long Short-term Memory (LSTM) network to learn aircraft trajectory patterns in the presence of various procedural constraints in airspace (Shi et al., 2020). Other research combined a Convolutional Neural Network (CNN) with an LSTM network to better extract spatio-temporal features in trajectories (Ma & Tian, 2020). Recently, various efforts have been made to apply Transformer-based architectures for trajectory prediction (Guo et al., 2022, 2024; Tong et al., 2023).
Due to the complex interactions between aircraft, air traffic controllers (ATCo) often manually assign specific headings, altitudes, and speeds to aircraft to ensure safe and efficient operation. This process is known as radar vectoring and results in diverse patterns in aircraft trajectories, as shown in Figure 1. However, ATCo can have particular preferences regarding how to deviate aircraft from predefined flight paths, which can lead to imbalance among different trajectory patterns. Figure 2 shows a histogram of different trajectory patterns for arriving aircraft at Inchon International Airport, South Korea, which illustrates the possibility of imbalance in the trajectory dataset. Such imbalance could deteriorate the performance of learning-based trajectory models and make them overfit major trajectory patterns. Therefore, it is beneficial to expand the dataset with synthetically generated data through dataset augmentation (Yoon & Lee, 2023).
Although a sufficient and balanced dataset is crucial for training data-driven aircraft trajectory models, very limited attention has been given to trajectory dataset augmentation. In this work, we propose a novel framework for aircraft trajectory dataset augmentation. The proposed framework consists of three parts. First, the original trajectory dataset is transformed into a learned latent space using a sequence autoencoder trained to reconstruct the input trajectory in its output. Various architectures, including a Sequence to Sequence (Seq2Seq) LSTM network (Sutskever et al., 2014), can be used as an autoencoder. However, we choose a Transformer-based architecture due to its superior representation capability for long sequences (Vaswani et al., 2017). Second, once the original trajectory dataset is transformed into a set of context vectors in the latent space, the probability distribution of the vectors is fitted by a GMM. Lastly, new samples are drawn from the fitted GMM and decoded in the same manner as in the trajectory reconstruction process in the first step.
@@
The main advantage of the proposed framework is that the underlying probability distribution of trajectory dataset in the latent space is explicitly estimated. By doing this, the newly augmented data can preserve the temporal dependency between trajectory points over time (i.e., aircraft dynamics) and the spatial flight patterns due to various operational constraints in specific airspace. The proposed framework for aircraft trajectory dataset augmentation is validated with actual air traffic surveillance data. Through a series of experiments, we demonstrate that the synthetic trajectory data generated by the proposed method outperform those by other generative models in both discriminative and predictive tasks.
@@ 2.1 Time-series data generation
Several studies have applied Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) to generate synthetic time-series data. One study used LSTM networks as both a generator and discriminator. A noise vector was used as input, and time-series data were generated recursively (Mogren, 2016). A similar approach was applied in another study, where more realistic time-series data were generated using a Recurrent Neural Network (RNN) with additional domain-specific inputs (Esteban et al., 2017). TimeGAN was proposed to better preserve the temporal dynamics in the time-series training data, and both reconstructive and adversarial objectives were jointly optimized (Yoon et al., 2019).
The Variational Autoencoders (VAEs) have been explored for time-series data generation. TimeVAE was introduced to enable user-defined distributions in the decoder and enrich the temporal properties of the generated data (Desai et al., 2021). The representation capabilities of transformers have led to their adoption in time-series generation tasks. For example, Museformer was designed with fine- and coarse-grained attention mechanisms for symbolic music generation (Yu et al., 2022). Another study combined an adversarial autoencoder with a Transformer to generate time-series data by integrating local and global features in the data (Srinivasan & Knottenbelt, 2022).
@@ 2.2 Aircraft trajectory generation
The generation of synthetic aircraft trajectories presents significant challenges because they need to capture both the temporal dynamics of aircraft movement and the spatial distribution while respecting various operational constraints within a specific airspace. For example, arriving aircraft should pass through specific three dimensional fixes and maintain appropriate altitudes to ensure safe and efficient landing. One such critical point is the Final Approach Fix (FAF), which marks the beginning of the final approach segment where pilots initiate their final descent for landing, as illustrated in Figure 3.
@@
One study applied a GMM to fit the distribution of an aircraft trajectory dataset and generate synthetic trajectories by drawing samples from the fitted GMM (Barratt et al., 2018). Another study also applied a GMM and used a feature vector that consisted of deviations from a flight path rather than actual position measurements to better represent variability in trajectories within specific airspace (Jung & Kochenderfer, 2023). The synthetic minority over-sampling technique (SMOTE) (Chawla et al., 2002) was employed to augment aircraft trajectory data in minor trajectory patterns to improve the accuracy of trajectory prediction (Yoon & Lee, 2023). Another study combined a Temporal Convolutional Network (TCN) and a VAE for aircraft trajectory generation (Krauth et al., 2023).
@@ 3.1.1 Self-attention mechanism
The self-attention mechanism is the heart of the Transformer (Vaswani et al., 2017) and mimics human cognitive processes by selectively focusing on specific elements within an input sequence. Attention mechanism is defined as:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ represent queries, keys, and values, respectively, and $d_{k}$ represents the dimension of the keys. The queries $\mathbf{Q}$ interact with the keys $\mathbf{K}$ through a dot-product operation and quantifies the relevance of each key to each query. This interaction produces attention weights that are normalized using the softmax function. These normalized weights are subsequently applied to the values $\mathbf{V}$, effectively contextualizing the input sequence. In practice, a multi-head attention is applied to provide a more diverse representation of the input sequence.
@@ 3.1.2 Gaussian mixture model
where $\mathbf{t}$ represents $C$-dimensional vector with a probability distribution comprising the weighted sum of $K$ Gaussian distributions. Here, $\pi_{i}$ denotes the mixture weight of the $i^{\text{th}}$ Gaussian component with a constraint of the sum of all mixture weights being equal to 1. The parameters $\{\pi_{i},\mu_{i},\Sigma_{i}\}$ of each $i^{\text{th}}$ Gaussian distribution are estimated using the Expectation-Maximization (EM) algorithm to maximize the likelihood of the observed data.
@@ 3.2 Proposed Framework
The proposed framework is illustrated in Figure 4. The framework consists of two main components: an autoencoder and a GMM. The autoencoder is trained to reconstruct the input trajectory and then used to convert the trajectory dataset from its original space into context vectors in the latent space. The decoder part of the autoencoder is used to synthesize new trajectory data from samples drawn from the GMM, which fits the probability distribution of the dataset in the latent space. Importantly, since the GMM training occurs after the autoencoder is fully trained, the generation results can be optimized without retraining the autoencoder by adjusting the associated hyperparameters (i.e., the number of principal components for PCA and the number of Gaussian distributions in the GMM).
@@ 3.2.1 Sequence Autoencoder
For the sequence autoencoder, we utilize a Transformer-based encoder with a self-attention mechanism and a Multi-Layer Perceptron (MLP) as a decoder. As illustrated in Figure 4(a), we first embed the aircraft trajectory $\{\mathbf{x}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times F\times S}$ into the embedding vectors with a hidden dimension $D$, where $B$ is the mini-batch size, $F$ is the number of original features (i.e., latitude, longitude, and altitude), and $S$ is the maximum length of aircraft trajectory sequences. The embedded trajectory sequence $\{\mathbf{x_{e}}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times D\times S}$ is then fed into the trajectory encoder.
The trajectory encoder employs the full self-attention mechanism, as proposed in the original Transformer model (Vaswani et al., 2017). This choice is based on long-term relationships between trajectory points at different timestamps being as crucial as their short-term relationships in air traffic operations. For example, the trajectory points in the early part of a sequence contain information about the entry point into the airspace and significantly influence the trajectory points at much later time steps due to the use of Standard Terminal Arrival Route (STAR). Finally, the latent vectors $\{\mathbf{z}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times D\times S}$ from the trajectory encoder are passed into the decoder, where the aircraft trajectory is reconstructed as $\{\mathbf{\tilde{x}}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times F\times S}$. We use L2 loss function, $\frac{1}{B}\Sigma_{i=1}^{B}\left\|\mathbf{x}^{i}-\mathbf{\tilde{x}}^{i}\right\|_{2}^{2}$, to measure the difference between the actual trajectories and the reconstructed trajectories. Notably, the use of the MLP in the trajectory decoder enables non-autoregressive trajectory generation. This significantly reduces the computation time and minimizes the risk of error accumulation, which typically occurs in other autoregressive times-series generative models.
@@ 3.2.2 Encoding and GMM
where $\overrightarrow{\mathbf{z}}^{j}\in\mathbb{R}^{1\times(D\cdot S)}$, and $\oplus$ denotes concatenation. We then stack $\overrightarrow{\mathbf{z}}^{j}$ row-wise, which results in a matrix $\mathbf{M}\in\mathbb{R}^{N\times(D\cdot S)}$. PCA is applied to this matrix, which results in a matrix $\mathbf{O}\in\mathbb{R}^{N\times P}$. Through PCA, the latent vectors $\overrightarrow{\mathbf{z}}^{j}$ are projected into a low-dimensional space $(1\leq P<(D\cdot S))$, where the principal feature vectors are orthogonal to each other. This step helps to manage redundancy and noise in the dataset and facilitates the training of the GMM, which fits the matrix $\mathbf{O}$ as a mixture of multiple Gaussian distributions $p(\mathbf{y})=\sum_{i=1}^{K}\pi_{i}\mathcal{N}(\mathbf{y};\mu_{i},\Sigma_{i})$.
@@ 3.2.3 Sampling and Generation
To augment a dataset, new latent vectors $\{\mathbf{y}^{q}\}^{M}_{q=1}\in\mathbb{R}^{M\times P}$ are sampled from the fitted GMM, where $M$ is the number of generated samples. They can then be reverted to the original dimension $(\mathbb{R}^{M\times P}\rightarrow\mathbb{R}^{M\times(D\cdot S)})$ and reshaped $(\mathbb{R}^{M\times(D\cdot S)}\rightarrow\mathbb{R}^{M\times D\times S})$ for decoding. Finally, new aircraft trajectory data $\{\tilde{\mathbf{y}}^{q}_{1:S}\}^{M}_{q=1}\in\mathbb{R}^{M\times F\times S}$ are generated as outputs of the decoder. Importantly, the new samples $\{\mathbf{y}^{q}\}^{M}_{q=1}$ can still be effectively decoded, as the multivariate Gaussian distribution $p(\mathbf{y})$ provides proper representations of the original dataset.
@@ 4 Numerical Experiments
The proposed framework was implemented in the following software and hardware environments: Python 3.9.16, PyTorch 2.0.1, CUDA 11.7, an Intel Core i7-12700F CPU, 32 GB of RAM, and an NVIDIA GeForce RTX 3060 GPU.
@@ 4.1 Dataset Description
We utilized trajectories of arriving aircraft in the terminal airspace (70-NM radius) at Incheon International Airport, South Korea. The dataset was acquired from FlightRadar24111https://www.flightradar24.com, which collects aircraft trajectory data via the Automatic Dependent Surveillance–Broadcast (ADS-B) system. The dataset covers the period from January 2022 to June 2022 and comprises a sequential record of the state of each aircraft. The data include details such as the time, aircraft type, geographical position (latitude, longitude, and altitude), and other pertinent information. The ADS-B trajectory data have a non-uniform sampling rate, so we resampled the data using a piecewise cubic Hermite interpolating polynomial (PCHIP) (Fritsch & Carlson, 1980) so that it would have a uniform rate ($\Delta t=6\ \text{sec}$). Table 1 provides a description of the ADS-B trajectory dataset used in this work.

### v3

@@ Abstract
Aircraft trajectory modeling plays a crucial role in air traffic management (ATM) and is important for various downstream tasks, including conflict detection and landing time prediction. Dataset augmentation by adding synthetically generated trajectory data is necessary to develop a more robust aircraft trajectory model and ensure that the trajectory dataset is sufficient and balanced. We propose a novel framework called ATRADA for aircraft trajectory dataset augmentation. In the proposed framework, a Transformer encoder learns the underlying patterns in the original trajectory dataset and converts each data point into a context vector in the learned latent space. The converted dataset is projected to reduced dimensions using principal component analysis (PCA), and a Gaussian mixture model (GMM) is applied to fit the probability distribution of the data points in the reduced-dimensional space. Finally, new samples are drawn from the fitted GMM, the dimension of the samples is reverted to the original dimension, and the samples are decoded with a multi-layer perceptron (MLP). Several experiments demonstrate that the framework effectively generates new, high-quality synthetic aircraft trajectory data, which were compared to the results of several baselines.
keywords:
Air traffic management, Aircraft trajectory dataset augmentation, Transformer, Gaussian mixture model
@@ 1 Introduction
As the demand for air transportation continues to increase around the world, the concept of air traffic operations has shifted to trajectory-based operation, which more accurately manages the four-dimensional trajectory of aircraft in terms of latitude, longitude, altitude, and time. Many studies have developed data-driven aircraft trajectory models for various downstream applications such as conflict detection and landing time prediction [1, 2].
One study suggested using a Gaussian mixture model (GMM) to learn statistical patterns of aircraft trajectories [3], while another study used a long short-term memory (LSTM) network to learn aircraft trajectory patterns in the presence of various procedural constraints in airspace [4]. Other research combined a convolutional neural network (CNN) with an LSTM network to better extract spatio-temporal features in trajectories [5]. Recently, various efforts have been made to apply Transformer-based architectures for trajectory prediction [6, 7, 8].
Due to the complex interactions between aircraft, air traffic controllers (ATCs) often manually assign specific headings, altitudes, and speeds to aircraft to ensure safe and efficient operation. This process is known as radar vectoring and results in diverse patterns in aircraft trajectories, as shown in Figure 1. However, ATCs can have particular preferences regarding how to deviate aircraft from predefined flight paths, which can lead to imbalance among different trajectory patterns. Figure 2 shows a histogram of different trajectory patterns for arriving aircraft at Inchon International Airport in South Korea, which illustrates the possibility of imbalance in the trajectory dataset. Such imbalance could deteriorate the performance of learning-based trajectory models and make them overfit major trajectory patterns. Therefore, it is beneficial to expand the dataset with synthetically generated data through dataset augmentation [9].
Although a sufficient and balanced dataset is important for training data-driven aircraft trajectory models, very limited attention has been given to trajectory dataset augmentation. We propose a novel framework for augmentation that features explicit estimation of the underlying probability distribution of a trajectory dataset in latent space. This allows the augmented data to preserve the temporal dependency between trajectory points over time (aircraft dynamics) and the spatial flight patterns necessitated by various operational constraints in a specific airspace. The framework was validated using actual air traffic surveillance data. A series of experiments demonstrates that the synthetic trajectory data generated by the proposed method outperform those of other generative models in both discriminative and predictive tasks.
@@ 2.1 Time-series data generation
Several studies have applied generative adversarial networks (GANs) [10] to generate synthetic time-series data. One study used LSTM networks as both a generator and discriminator. A noise vector was used as input, and time-series data were generated recursively [11]. Another study applied a similar approach and generated more realistic time-series data using a recurrent neural network (RNN) with additional domain-specific inputs [12]. TimeGAN was proposed to better preserve the temporal dynamics in the time-series training data, and both reconstructive and adversarial objectives were jointly optimized [13].
Variational autoencoders (VAEs) have been explored for time-series data generation. TimeVAE was introduced to enable user-defined distributions in the decoder and enrich the temporal properties of the generated data [14]. The representation capabilities of transformers have led to their adoption in time-series generation tasks. For example, Museformer was designed with fine- and coarse-grained attention mechanisms for symbolic music generation [15]. Another study combined an adversarial autoencoder with a Transformer to generate time-series data by integrating local and global features in the data [16].
@@ 2.2 Aircraft trajectory generation
The generation of synthetic aircraft trajectories presents significant challenges because the trajectories need to capture both the temporal dynamics of aircraft movement and the spatial distribution while respecting various operational constraints within a specific airspace. For example, arriving aircraft must pass through specific three-dimensional fixes and maintain appropriate altitudes to ensure safe and efficient landing. One such critical point is the final approach fix (FAF), which marks the beginning of the final approach segment where pilots initiate their final descent for landing, as illustrated in Figure 3.
@@
One study applied a GMM to fit the distribution of an aircraft trajectory dataset and generate synthetic trajectories by drawing samples from the fitted GMM [3]. Another study also applied a GMM and used a feature vector that consisted of deviations from a flight path rather than actual position measurements to better represent variability in trajectories within specific airspace [17]. The synthetic minority over-sampling technique (SMOTE) [18] was used to augment aircraft trajectory data in minor trajectory patterns to improve the accuracy of trajectory prediction [9]. Another study combined a temporal convolutional network (TCN) and a VAE for aircraft trajectory generation [19].
@@ 3 Methodology
The proposed framework consists of three parts. First, the original trajectory dataset is transformed into a learned latent space using a sequence autoencoder trained to reconstruct the input trajectory in its output. Various architectures, including a sequence to sequence (Seq2Seq) LSTM network [20], can be used as an autoencoder. However, we chose a Transformer-based architecture due to its better representation capability for long sequences [21]. Second, once the original trajectory dataset is transformed into a set of context vectors in the latent space, the probability distribution of the vectors is fitted by a GMM. Lastly, new samples are drawn from the fitted GMM and decoded in the same manner as in the trajectory reconstruction process in the first step.
@@ 3.1.1 Self-attention mechanism
The self-attention mechanism is the heart of the Transformer [21] and mimics human cognitive processes by selectively focusing on specific elements within an input sequence. The attention mechanism is defined as:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ represent queries, keys, and values, respectively, and $d_{k}$ represents the dimension of the keys. The queries $\mathbf{Q}$ interact with the keys $\mathbf{K}$ through a dot-product operation that quantifies the relevance of each key to each query. This interaction produces attention weights that are normalized using the softmax function, which are applied to the values $\mathbf{V}$ to contextualize the input sequence. In practice, multi-head attention is applied to provide a more diverse representation of the input sequence.
@@ 3.1.2 Gaussian mixture model
where $\mathbf{t}$ is a $C$-dimensional vector with a probability distribution comprising the weighted sum of $K$ Gaussian distributions. $\pi_{i}$ denotes the mixture weight of the $i^{\text{th}}$ Gaussian component, which has a constraint that the sum of all mixture weights must be equal to 1. The parameters $\{\pi_{i},\mu_{i},\Sigma_{i}\}$ of each $i^{\text{th}}$ Gaussian distribution are estimated using the expectation-maximization (EM) algorithm to maximize the likelihood of the observed data.
@@ 3.2 Proposed framework
As illustrated in Figure 4, the proposed framework consists of two main components: an autoencoder and a GMM. The autoencoder is trained to reconstruct the input trajectory and then used to convert the trajectory dataset from its original space into context vectors in the latent space. The decoder part of the autoencoder synthesizes new trajectory data from samples drawn from the GMM, which fits the probability distribution of the dataset in the latent space. Importantly, since the GMM training occurs after the autoencoder is fully trained, the generation results can be optimized without retraining the autoencoder by adjusting the associated hyperparameters: the number of principal components for principal component analysis (PCA) and the number of Gaussian distributions in the GMM.
@@ 3.2.1 Sequence autoencoder
For the sequence autoencoder, we use a Transformer-based encoder with a self-attention mechanism and a multi-layer perceptron (MLP) as a decoder. As illustrated in Figure 4(a), we first embed the aircraft trajectory $\{\mathbf{x}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times F\times S}$ into the embedding vectors with a hidden dimension $D$, where $B$ is the mini-batch size, $F$ is the number of original features (latitude, longitude, and altitude), and $S$ is the maximum length of aircraft trajectory sequences. The embedded trajectory sequence $\{\mathbf{x_{e}}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times D\times S}$ is then fed into the trajectory encoder.
The trajectory encoder uses the full self-attention mechanism, as proposed in the original Transformer model [21]. This choice is based on the equal importance of the long-term relationships between trajectory points at different timestamps to that of the short-term relationships in air traffic operations. For example, the trajectory points in the early part of a sequence contain information about the entry point into the airspace and significantly influence the trajectory points at much later time steps due to the use of standard terminal arrival route (STAR). Finally, the latent vectors $\{\mathbf{z}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times D\times S}$ from the trajectory encoder are passed into the decoder, where the aircraft trajectory is reconstructed as $\{\mathbf{\tilde{x}}^{i}_{1:S}\}^{B}_{i=1}\in\mathbb{R}^{B\times F\times S}$. We measure the difference between the actual trajectories and the reconstructed trajectories using an L2 loss function, $\frac{1}{B}\Sigma_{i=1}^{B}\left\|\mathbf{x}^{i}-\mathbf{\tilde{x}}^{i}\right\|_{2}^{2}$. Notably, the use of the MLP in the trajectory decoder enables non-autoregressive trajectory generation, which significantly reduces the computation time and minimizes the risk of error accumulation that typically occurs in other autoregressive times-series generative models.
@@ 3.2.2 Encoding and GMM
where $\overrightarrow{\mathbf{z}}^{j}\in\mathbb{R}^{1\times(D\cdot S)}$, and $\oplus$ denotes concatenation. $\overrightarrow{\mathbf{z}}^{j}$ is then stacked row-wise, which results in a matrix $\mathbf{M}\in\mathbb{R}^{N\times(D\cdot S)}$. PCA is applied to this matrix, which results in a matrix $\mathbf{O}\in\mathbb{R}^{N\times P}$. Through PCA, the latent vectors $\overrightarrow{\mathbf{z}}^{j}$ are projected into a low-dimensional space $(1\leq P<(D\cdot S))$, where the principal feature vectors are orthogonal to each other. This step helps to manage redundancy and noise in the dataset and facilitates the training of the GMM, which fits the matrix $\mathbf{O}$ as a mixture of multiple Gaussian distributions $p(\mathbf{y})=\sum_{i=1}^{K}\pi_{i}\mathcal{N}(\mathbf{y};\mu_{i},\Sigma_{i})$.
@@ 3.2.3 Sampling and generation
To augment a dataset, new latent vectors $\{\mathbf{y}^{q}\}^{M}_{q=1}\in\mathbb{R}^{M\times P}$ are sampled from the fitted GMM, where $M$ is the number of generated samples. The vectors can then be reverted to the original dimension $(\mathbb{R}^{M\times P}\rightarrow\mathbb{R}^{M\times(D\cdot S)})$ and reshaped $(\mathbb{R}^{M\times(D\cdot S)}\rightarrow\mathbb{R}^{M\times D\times S})$ for decoding. New aircraft trajectory data $\{\tilde{\mathbf{y}}^{q}_{1:S}\}^{M}_{q=1}\in\mathbb{R}^{M\times F\times S}$ are then generated as outputs of the decoder. Importantly, the new samples $\{\mathbf{y}^{q}\}^{M}_{q=1}$ can still be effectively decoded, as the multivariate Gaussian distribution $p(\mathbf{y})$ provides proper representations of the original dataset.
@@ 4 Numerical Experiments
The proposed framework was implemented with the following software and hardware: Python 3.9.16, PyTorch 2.0.1, CUDA 11.7, an Intel Core i7-12700F CPU, 32 GB of RAM, and an NVIDIA GeForce RTX 3060 GPU.
@@ 4.1 Dataset description
We utilized trajectories of arriving aircraft in the terminal airspace (70-NM radius) at Incheon International Airport. The dataset was acquired from FlightRadar24111https://www.flightradar24.com, which collects aircraft trajectory data via the automatic dependent surveillance–broadcast (ADS-B) system. The dataset covers the period of January 2022 to June 2022 and comprises sequential records of the state of each aircraft. The data include details such as the time, aircraft type, geographical position (latitude, longitude, and altitude), and other pertinent information. The ADS-B trajectory data have a non-uniform sampling rate, so we resampled the data using a piecewise cubic Hermite interpolating polynomial (PCHIP) [22] to make the rate uniform ($\Delta t=6\ \text{sec}$). Table 1 provides a description of the ADS-B trajectory dataset used in this work.